# The Effectiveness of Large Language Models in Transforming Unstructured Text to Standardized Formats

### v1

@@ Abstract
The exponential growth of unstructured text data presents a fundamental challenge in modern data management and information retrieval. While Large Language Models (LLMs) have shown remarkable capabilities in natural language processing, their potential to transform unstructured text into standardized, structured formats remains largely unexplored - a capability that could revolutionize data processing workflows across industries. This study breaks new ground by systematically evaluating LLMs’ ability to convert unstructured recipe text into the structured Cooklang format. Through comprehensive testing of four models (GPT-4o, GPT-4o-mini, Llama3.1:70b, and Llama3.1:8b), an innovative evaluation approach is introduced that combines traditional metrics (WER, ROUGE-L, TER) with specialized metrics for semantic element identification. Our experiments reveal that GPT-4o with few-shot prompting achieves breakthrough performance (ROUGE-L: 0.9722, WER: 0.0730), demonstrating for the first time that LLMs can reliably transform domain-specific unstructured text into structured formats without extensive training. Although model performance generally scales with size, we uncover surprising potential in smaller models like Llama3.1:8b for optimization through targeted fine-tuning. These findings open new possibilities for automated structured data generation across various domains, from medical records to technical documentation, potentially transforming the way organizations process and utilize unstructured information.
@@ 1 Introduction
The exponential growth of unstructured text data presents a challenge in data management, analysis, and information retrieval. While Large Language Models (LLMs)Brown et al. (2020) have demonstrated remarkable capabilities in natural language processing, their potential to transform unstructured text into standardized, structured domain-specific formats remains unexplored. This capability could revolutionize data processing workflows across various domains, enabling more efficient information retrieval, automated analysis, and data management.
@@
In order to evaluate the performance of LLM systems in the generation of structured recipes, a variety of methods are employed for both qualitative and quantitative assessment. Our methodology included traditional natural language processing metrics such as Word Error Rate (WER) Ali and Renals (2018); Klakow and Peters (2002), and ROUGE-L Research (2019); See et al. (2017) which measure the accuracy and fluency of the generated text. Additionally, we introduce domain-specific metrics to evaluate the accuracy of a generated recipe. These include an Ingredient Identification Score and Unit and Amount Identification Scores. The evaluation of these metrics allows for the assessment of not only the linguistic quality of the generated recipes but also their adherence to the specific structural requirements of the Cooklang format and their culinary accuracy. Furthermore, we analyze the LLMs performance across different input configurations and prompting techniques (zero-shot PromptingGuide (2023a), few-shot PromptingGuide (2023b), and MIPROv2 Opsahl-Ong et al. (2024)). This evaluation approach enables us to understand the strengths and limitations of various LLMs in the context of structured domain-specific format generation.
@@ 2 Background
One of the use cases for LLMs is the extraction of food entities from cooking recipes, as demonstrated in the study by Pitsilou et al. Pitsilou et al. (2024). The results of this study demonstrate that, even in the absence of labeled training data, LLMs can achieve promising results in domain-specific tasks such as named entity recognition (NER) for food items. Another illustration of the application of LLMs for structured text parsing outside the culinary domain is StructGPT, a general framework to enhance LLMs zero-shot reasoning capabilities over structured data such as knowledge graphs, tables, and databases Jiang et al. (2023). An additional example of the use of LLMs in a culinary context is LLaVA-Chef, a multi-modal generative model specifically designed for food recipes Mohbat and Zaki (2024). By adapting the LLaVA (Large Language and Vision Assistant) Liu et al. (2024) model to the food domain through a multi-stage fine-tuning approach, the authors achieved state-of-the-art performance in recipe generation tasks. Notably, these results demonstrate competitive performance in comparison to the work of Patil et al. Pansare et al. (2024), who employed convolutional neural networks (CNNs) for image recognition in conjunction with deep learning models for natural language processing. Another illustration of the application of LLMs to culinary tasks is the development of RecipeNLG Bień et al. (2020), which employed language models to enhance semi-structured recipe text generation and generate a dataset for subsequent LLM training and refinement. As this domain progresses, it demonstrates the potential for integrating general-purpose language models with specialized domain expertise. Building on this work, Zhou et al.’s research on FoodSky Zhou et al. (2024) demonstrated that LLMs can achieve better performance in domain-specific tasks by using specialized mechanisms like the Topic-based Selective State Space Model (TS3M) and Hierarchical Topic Retrieval Augmented Generation (HTRAG). The study demonstrated the effectiveness of FoodSky by achieving accuracy rates of 67.2% and 66.4% in chef and dietetic examinations, respectively. Authors suggest that LLMs, when adequately trained and enhanced with domain-specific knowledge, can efficiently handle structured text conversion tasks. This lays a solid foundation for applications in recipe processing and other specialized fields. The research paper Cook2LTL Mavrogiannis et al. (2024) demonstrates how to utilize a language model to translate recipes into linear temporal logic (LTL). Cook2LTL shows how structured representations can bridge the gap between human-readable content and machine-executable instructions. Another application of LLM Dubovskoy (2024) employs graph-based representations in order to track ingredient transformations and state changes throughout the cooking process. This evolution, from basic formalization to pattern-based approaches and finally to LLM-powered graph representations, illustrates the ongoing efforts to bridge the gap between natural language cooking instructions and structured, computationally accessible formats.
@@
As LLMs have gained popularity, prompt engineering Brown et al. (2020) has emerged as a compelling alternative to fine-tuning approaches. Research of language models has focused on developing and evaluating prompt techniques across a diverse range of tasks. The research space has explored multiple prompting approaches: zero-shot learning PromptingGuide (2023a), where models perform tasks without prior examples; few-shot learning PromptingGuide (2023b), which utilizes a small number of examples to guide performance, or more complex prompt techniques like MIPRO Opsahl-Ong et al. (2024) which optimizes both instructions and few-shot demonstrations for multi-stage language model programs. The evolution of prompting techniques has demonstrated key advantages over traditional fine-tuning methods.
@@
Prompt engineering requires less computational resources and training data compared to full model fine-tuning. The well-built prompts can generate strong performance Li et al. (2023) across a range of tasks while maintaining model flexibility. These prompts have shown remarkable effectiveness across applications: in sentiment analysis, prompt-based fine-tuning achieved 92.7% accuracy compared to traditional fine-tuning’s 81.4% while requiring only 32 examples on the SST-2 dataset Gao et al. (2021). Similarly, in classification tasks Wei et al. (2024) on datasets like SNLI, prompt-based methods with demonstrations achieved 79.7% accuracy compared to standard fine-tuning’s 48.4%, demonstrating the approach’s effectiveness across different linguistic tasks. Furthermore, prompting techniques can enhance the reasoning capabilities of language models. A notable example is the self-consistency approach introduced by Wang et al. Wang et al. (2022), which builds upon chain-of-thought prompting to improve complex reasoning tasks. This method leverages the intuition that multiple reasoning paths can lead to a correct answer, similar to human problem-solving processes. By sampling diverse reasoning paths instead of relying on a single greedy decoding path, their approach achieved striking improvements across reasoning benchmarks: GSM8K, SVAMP, and AQuA.
@@ 3 Methodology
These LLMs represent the most popular models from both closed and open source, allowing for an assessment of performance across a range of capabilities. Through the testing of combinations of input formats, prompting strategies, and model architectures, the objective was to identify the most effective approach for converting recipe texts into the structured domain-specific recipe format. The recipe format contains two distinct input fields, as shown in dspy.Signatures (Listings LABEL:lst:dspy_text2cooklang, LABEL:lst:dspy_text2cooklang_without_cooklang, and LABEL:lst:dspy_text2cooklang_without_cooklang_without_ings). The first field contains ingredients for the recipe, presented as free text where each ingredient is comma-separated. The second field consists of the general recipe text, primarily this field comprises cooking instructions. This field is a converted Cooklang format via LLM. The objective of this research is to examine the limitations of LLM in structured domain-specific text generation. The research explores how different models and prompt engineering influence semantic preservation and structural consistency in the conversion of unstructured recipes to formal specifications. Through an evaluation of model size and prompt engineering approaches, the research aims to advance our understanding of LLMs capabilities in maintaining semantic fidelity while adhering to strict syntactic constraints. This fundamental challenge in natural language processing extends far beyond recipe conversion to broader applications in structured knowledge representation.
@@ 3.1 Prompt engineering
We based our approach on transforming recipes into Cooklang format using an LLM on the DSPy framework Khattab et al. (2024). We developed distinct DSPy programs for each model input type to ensure optimal performance across all scenarios. We implemented few-shot prompting using the bootstrap-few-shot-with-random-search method from the DSPy library, keeping the default hyperparameters. Similarly, we implemented MIPROv2 Opsahl-Ong et al. (2024) using the existing DSPy implementation with its default hyperparameters. The Listings LABEL:lst:dspy_text2cooklang, LABEL:lst:dspy_text2cooklang_without_cooklang show the DSPy programs we used in our implementation. These prompts form the core elements of our recipe transformation process.

### v2

@@ Abstract
The exponential growth of unstructured text data presents a fundamental challenge in modern data management and information retrieval. While Large Language Models (LLMs) have shown remarkable capabilities in natural language processing, their potential to transform unstructured text into standardized, structured formats remains largely unexplored - a capability that could revolutionize data processing workflows across industries. This study breaks new ground by systematically evaluating LLMs’ ability to convert unstructured recipe text into the structured Cooklang format. Through comprehensive testing of four models (GPT-4o, GPT-4o-mini, Llama3.3:70b, and Llama3.1:8b), an innovative evaluation approach is introduced that combines traditional metrics (WER, ROUGE-L, TER) with specialized metrics for semantic element identification. Our experiments reveal that GPT-4o with few-shot prompting achieves breakthrough performance (ROUGE-L: 0.8209, WER: 0.3509), demonstrating for the first time that LLMs can reliably transform domain-specific unstructured text into structured formats without extensive training. Although model performance generally scales with size, we uncover surprising potential in smaller models like Llama3.1:8b for optimization through targeted fine-tuning. These findings open new possibilities for automated structured data generation across various domains, from medical records to technical documentation, potentially transforming the way organizations process and utilize unstructured information.
@@ 1 Introduction
The exponential growth of unstructured text data presents a challenge in data management, analysis, and information retrieval. While Large Language Models (LLMs) Brown and et al. (2020) have demonstrated remarkable capabilities in natural language processing, their potential to transform unstructured text into standardized, structured domain-specific formats remains unexplored. This capability could revolutionize data processing workflows across various domains, enabling more efficient information retrieval, automated analysis, and data management.
@@
In order to evaluate the performance of LLM systems in the generation of structured recipes, a variety of methods are employed for both qualitative and quantitative assessment. Our methodology included traditional natural language processing metrics such as Word Error Rate (WER) Ali and Renals (2018); Klakow and Peters (2002), and ROUGE-L Research (2019); See et al. (2017), which measure the accuracy and fluency of the generated text. Additionally, we introduce domain-specific metrics to evaluate the accuracy of a generated recipe. These include an Ingredient Identification Score and Unit and Amount Identification Scores. The evaluation of these metrics allows for the assessment of not only the linguistic quality of the generated recipes but also their adherence to the specific structural requirements of the Cooklang format and their culinary accuracy. Furthermore, we analyze the LLMs performance across different input configurations and prompting techniques (zero-shot PromptingGuide (2023a), few-shot PromptingGuide (2023b), and MIPROv2 Opsahl-Ong et al. (2024)). This evaluation approach enables us to understand the strengths and limitations of various LLMs in the context of structured domain-specific format generation.
In this study, we aim to address two research questions: How do different LLM sizes affect the quality and accuracy of structured recipe generation? What impact do various prompting techniques have on the generation of well-formed Cooklang specifications? The applications of this research extend beyond the culinary domain, with potential implications for healthcare documentation using HL7 format, music notation in MusicXML, and technical documentation organization. The key contributions of this research include:
  • A systematic evaluation of LLMs’ ability to transform unstructured recipe text into structured Cooklang format across different models and sizes.
  • Identification of scaling patterns across model sizes, with the surprising finding that smaller models like Llama3.1:8b show potential for optimization through targeted fine-tuning.
  • Demonstration that LLMs can reliably transform domain-specific unstructured text into structured formats without requiring extensive training.
  • Empirical evidence that GPT-4o with few-shot prompting achieves breakthrough performance (ROUGE-L: 0.8209, WER: 0.3509) in structured format conversion.
  • A framework for implementing domain-specific evaluation methodology that combines traditional metrics (WER, ROUGE-L, TER) with specialized metrics for semantic element identification in recipes.
  • Insights that open new possibilities for automated structured data generation across various domains, including healthcare, technical documentation, and legal contracts
@@ 2 Background
One of the use cases for LLMs is the extraction of food entities from cooking recipes, as demonstrated in the study by Pitsilou et al. Pitsilou et al. (2024). The results of this study demonstrate that, even in the absence of labeled training data, LLMs can achieve promising results in domain-specific tasks such as named entity recognition (NER) for food items. Another illustration of the application of LLMs for structured text parsing outside the culinary domain is StructGPT, a general framework to enhance LLMs’ zero-shot reasoning capabilities over structured data such as knowledge graphs, tables, and databases Jiang et al. (2023). An additional example of the use of LLMs in a culinary context is LLaVA-Chef, a multi-modal generative model specifically designed for food recipes Mohbat and Zaki (2024). By adapting the LLaVA (Large Language and Vision Assistant) Liu et al. (2024) model to the food domain through a multi-stage fine-tuning approach, the authors achieved state-of-the-art performance in recipe generation tasks. Notably, these results demonstrate competitive performance in comparison to the work of Patil et al. Pansare et al. (2024), who employed convolutional neural networks (CNNs) for image recognition in conjunction with deep learning models for natural language processing. Another illustration of the application of LLMs to culinary tasks is the development of RecipeNLG Bień et al. (2020), which employed language models to enhance semi-structured recipe text generation and generate a dataset for subsequent LLM training and refinement. As this domain progresses, it demonstrates the potential for integrating general-purpose language models with specialized domain expertise. Building on this work, Zhou et al.’s research on FoodSky Zhou et al. (2024) demonstrated that LLMs can achieve better performance in domain-specific tasks by using specialized mechanisms like the Topic-based Selective State Space Model (TS3M) and Hierarchical Topic Retrieval Augmented Generation (HTRAG). The study demonstrated the effectiveness of FoodSky by achieving accuracy rates of 67.2% and 66.4% in chef and dietetic examinations, respectively. Authors suggest that LLMs, when adequately trained and enhanced with domain-specific knowledge, can efficiently handle structured text conversion tasks. This lays a solid foundation for applications in recipe processing and other specialized fields. The research paper Cook2LTL Mavrogiannis et al. (2024) demonstrates how to utilize a language model to translate recipes into linear temporal logic (LTL). Cook2LTL shows how structured representations can bridge the gap between human-readable content and machine-executable instructions. Another application of LLM Dubovskoy (2024) employs graph-based representations in order to track ingredient transformations and state changes throughout the cooking process. This evolution, from basic formalization to pattern-based approaches and finally to LLM-powered graph representations, illustrates the ongoing efforts to bridge the gap between natural language cooking instructions and structured, computationally accessible formats.
Study "Fine-tuning Language Models for Recipe Generation: A Comparative Analysis and Benchmark Study" Vij et al. (2025) investigated the performance of smaller language models, including T5-small, SmolLM, and Phi-2, specifically for recipe generation tasks. The authors evaluated these models using both traditional NLP metrics (BLEU, ROUGE) and domain-specific metrics such as Ingredient Coverage and Step Complexity. While this study did not include larger models like GPT-4o or Llama3.1:8b that are examined in our work, it provides an important benchmark within the recipe domain that helps contextualize our findings. For broader model comparisons, established leaderboards such as the Hugging Face Open LLM Leaderboard222hf.com/open_llm_leaderboard and LM Arena [^3] https://lmarena.ai/ offer comprehensive evaluations across various benchmarks, which further support our experimental design and analysis framework.
@@
As LLMs have gained popularity, prompt engineering Brown and et al. (2020) has emerged as a compelling alternative to fine-tuning approaches. Research on language models has focused on developing and evaluating prompt techniques across a diverse range of tasks. The research space has explored multiple prompting approaches: zero-shot learning PromptingGuide (2023a), where models perform tasks without prior examples; few-shot learning PromptingGuide (2023b), which utilizes a small number of examples to guide performance, or more complex prompt techniques like MIPRO Opsahl-Ong et al. (2024), which optimizes both instructions and few-shot demonstrations for multi-stage language model programs. The evolution of prompting techniques has demonstrated key advantages over traditional fine-tuning methods.
@@
Prompt engineering requires fewer computational resources and training data compared to full model fine-tuning. The well-built prompts can generate strong performance Li et al. (2023) across a range of tasks while maintaining model flexibility. These prompts have shown remarkable effectiveness across applications: in sentiment analysis, prompt-based fine-tuning achieved 92.7% accuracy compared to traditional fine-tuning’s 81.4% while requiring only 32 examples on the SST-2 dataset Gao et al. (2021). Similarly, in classification tasks Wei et al. (2024) on datasets like SNLI, prompt-based methods with demonstrations achieved 79.7% accuracy compared to standard fine-tuning’s 48.4%, demonstrating the approach’s effectiveness across different linguistic tasks. Furthermore, prompting techniques can enhance the reasoning capabilities of language models. A notable example is the self-consistency approach introduced by Wang et al. Wang et al. (2022), which builds upon chain-of-thought prompting to improve complex reasoning tasks. This method leverages the intuition that multiple reasoning paths can lead to a correct answer, similar to human problem-solving processes. By sampling diverse reasoning paths instead of relying on a single greedy decoding path, their approach achieved striking improvements across reasoning benchmarks: GSM8K, SVAMP, and AQuA.
@@
While existing research demonstrates the significant progress in LLM applications for domain-specific tasks and the evolution of various prompting techniques, a critical gap remains in systematically evaluating LLMs’ capability to transform unstructured text into standardized structured formats across different model architectures and prompting strategies. Prior studies have primarily focused on using LLMs for domain-specific knowledge extraction (as seen in food entity extraction research) or generation tasks (exemplified by RecipeNLG and LLaVA-Chef) but have not comprehensively assessed how different LLM architectures and prompting techniques affect the quality and accuracy of format conversion tasks. This gap is particularly significant as it represents the intersection of natural language understanding, structured data generation, and domain-specific knowledge representation—a capability with broad applications beyond culinary contexts to fields such as healthcare documentation, technical specification authoring, and legal document processing. Our work addresses this gap by providing a systematic comparative analysis of how different LLMs models sizes (from 8b to 70b parameters for open source) and prompting techniques (zero-shot, few-shot, and MIPROv2) influence performance in transforming unstructured recipe text into the standardized Cooklang format, offering insights applicable to structured text conversion challenges across domains.
@@ 3 Methodology
These LLMs represent the most popular models from both closed and open source, allowing for an assessment of performance across a range of capabilities. Through the testing of combinations of input formats, prompting strategies, and model architectures, the objective was to identify the most effective approach for converting recipe texts into the structured domain-specific recipe format. The recipe format contains two distinct input fields, as shown in dspy.Signatures (Listings LABEL:lst:dspy_text2cooklang, LABEL:lst:dspy_text2cooklang_without_cooklang, and LABEL:lst:dspy_text2cooklang_without_cooklang_without_ings). The first field contains ingredients for the recipe, presented as free text where each ingredient is comma-separated. The second field consists of the general recipe text, primarily, this field comprises cooking instructions. This field is converted to Cooklang format via LLM. The objective of this research is to examine the limitations of LLM in structured domain-specific text generation. The research explores how different models and prompt engineering influence semantic preservation and structural consistency in the conversion of unstructured recipes to formal specifications. Through an evaluation of model size and prompt engineering approaches, the research aims to advance our understanding of LLMs capabilities in maintaining semantic fidelity while adhering to strict syntactic constraints. This fundamental challenge in natural language processing extends far beyond recipe conversion to broader applications in structured knowledge representation.
Recipe formatting serves as an ideal benchmark for structured text conversion for several key reasons:
  • Dual nature of content: Recipes naturally combine narrative instructions with structured data elements (ingredients, measurements, cooking times), making them an excellent test case for evaluating both semantic understanding and structured data extraction.
  • Entity complexity: The Cooklang format requires accurate identification of multiple entity types (ingredients, cookware, timers, comments) with their associated attributes (amounts, units), testing an LLM’s fine-grained entity recognition capabilities.
  • Balancing precision and semantics: Successful conversion requires both preserving the semantic meaning of instructions while adhering to strict syntactic rules defined by the Cooklang specification.
  • Domain transferability: The challenges in recipe formatting—identifying entities, preserving relationships, maintaining semantic equivalence—parallel those in other domains such as healthcare documentation, technical specifications, and legal documents.
  • Clear evaluation criteria: Recipe formatting provides objectively measurable outcomes (correct ingredient identification, proper syntax application) that allow for quantitative assessment of model performance.
  • Accessibility: Unlike highly specialized domains, recipes are universally understood, making the benchmark accessible while still maintaining sufficient complexity to challenge state-of-the-art models.
@@ 3.1 Prompt engineering
We based our approach on transforming recipes into Cooklang format using an LLM on the DSPy framework Khattab et al. (2024). We developed distinct DSPy programs for each model input type to ensure optimal performance across all scenarios. We implemented few-shot prompting using the bootstrap-few-shot-with-random-search method from the DSPy library, keeping the default hyperparameters. Similarly, we implemented MIPROv2 Opsahl-Ong et al. (2024) using the existing DSPy implementation with its default hyperparameters. The Listings LABEL:lst:dspy_text2cooklang, LABEL:lst:dspy_text2cooklang_without_cooklang show the DSPy programs we used in our implementation. These prompts form the core elements of our recipe transformation process. Examples of prompts for Few-Shot and MIPROv2 could be found in Listings LABEL:fig:cooklang-prompt-1 and LABEL:fig:cooklang-prompt-2, respectively.
1Convert plain recipe text with provided ingredients into Cooklang text format.
2Return only Cooklang formatted recipe, dont return any other information. Return whole recipe in Cooklang format! Dont stop till you reach the end of the recipe.
3
4---
5
6Ingredients: 50 g ciabatta,some oil,,50 g cottage cheese,80 g smoked salmon,5 g capers,1 g dill,some lemon,some black pepper
7Recipe Text: Slice the whole ciabatta loaf in half and then slice the halves lengthwise. Fry the ciabatta halves in a frying pan in a little oil, a griddle pan gives the best results and makes nice lines as the bread toasts. Spread the cottage cheese on the toasted ciabatta liberally, cut out the smoked salmon slices and place on the bread with a sprinkling of capers and finally chopped dill. Serve with a lemon wedge and freshly cracked black pepper.
8Cooklang: Slice the whole @ciabatta{50%g} loaf in half and then slice the halves lengthwise. Fry the ciabatta halves in a #frying pan{} in a little @oil, a griddle pan gives the best results and makes nice lines as the bread toasts. Spread the @cottage cheese{50%g} on the toasted ciabatta liberally, cut out the @smoked salmon{80%g} slices and place on the bread with a sprinkling of @capers{5%g} and finally chopped @dill{1%g}. Serve with a @lemon wedge and freshly cracked @black pepper{}.
9
10[Additional examples omitted for brevity]
11
12---
13
14Follow the following format.
15
16Ingredients: Ingredients for the recipe. Comma separated list of ingredients.
17
18Recipe Text: Recipe text to convert to Cooklang format.
19
20Reasoning: Lets think step by step in order to \${produce the cooklang}. We ...
21
22Cooklang: Cooklang formatted recipe.
23
24---
25
26Ingredients: {ingredients}
27
28Recipe Text: {recipe text}
29
30Reasoning: Lets think step by step in order to Cooklang:
Listing 4: Example of prompt (Few-Shot) for Cooklang conversion without cooklang notation and with ingredients
1Transform the given plain text recipe and accompanying ingredients into a structured Cooklang format. Use the following specifications:
2
31. **Ingredients:**
4 - Prefix each ingredient with @‘.
5 - For multi-word ingredients, wrap the entire ingredient in ‘{}‘.
6 - Indicate the quantity directly after the ingredient name in ‘{}‘, followed by a % for the unit.
7 - Example: @salt‘, @ground black pepper{}‘, @potato{2}‘, @bacon strips{1%kg}‘, @syrup{1/2%tbsp}‘.
8
92. **Comments:**
10 - Use ‘--‘ to denote single-line comments.
11 - For multi-line comments, enclose them in ‘[- -]‘.
12 - Example: ‘-- Dont burn the roux!‘.
13
143. **Cookware:**
15 - Use # to specify the cookware required.
16 - For multi-word items, enclose in ‘{}‘.
17 - Example: #pot‘, #potato masher{}‘.
18
194. **Timers:**
20 - Use ‘~‘ to specify timing instructions.
21 - Indicate the duration in ‘{}‘.
22 - You may include a name before the duration.
23 - Example: ‘~{25%minutes}‘, ‘~eggs{3%minutes}‘.
24
25Return only the Cooklang formatted recipe without any additional commentary or information. Ensure that the output is complete and captures all elements of the recipe succinctly.
26
27---
28
29[Few-shot examples omitted for brevity]
30
31---
32
33Follow the following format.
34
35Ingredients: Ingredients for the recipe. Comma-separated list of ingredients.
36
37Recipe Text: Recipe text to convert to Cooklang format.
38
39Reasoning: Lets think step by step in order to \${produce the cooklang}. We ...
40
41Cooklang: Cooklang formatted recipe.
42
43---
44
45Ingredients: {ingredients}
46
47Recipe Text: {recipe text}
48
49Reasoning: Lets think step by step in order to
Listing 5: Example of prompt (MIPROv2) for Cooklang conversion with cooklang notation and with ingredients